# Short-distance thermal phase structure of charged black holes in 4D Einstein–Gauss–Bonnet gravity

### v4

@@ V.2 Computing the Ruppeiner thermodynamic curvature $R_{C}$
where $g=\det{g_{\mu\nu}}=g_{MM}g_{QQ}-g_{MQ}^{2}$.
In this framework, the following interpretation is usually ascribed to $R_{C}$: a zero curvature indicates that the system is non-interacting and in an ideal gas-like configuration, providing no additional information about black hole micromolecules. A positive $R$ is associated with repulsive effective interactions, while negative $R$ indicates attractive interactions. Divergences in $R$ correspond to phase transitions Ruppeiner (2014). From this interpretation of $R$, one might wonder if $R$ would be very large for a black hole due to its infinite density. This point is emphasized in Ref. Ruppeiner (2014), where it is suggested that the gravitational degrees of freedom in a black hole system might possess some non-statistical description because all the gravitating material has been compressed into the central singularity. Thus, thermodynamic curvature represents some kind of non-gravitational interactions among the black hole constituents at its surface, arising effectively from the underlying gravitational degrees of freedom.
Figure 6: Thermodynamic curvature $R_{C}$ vs. black hole size: impact of (a) quantum corrections $\eta$, (b) positive GB coupling parameter $\alpha$, and (c) negative GB coupling parameter $\alpha$; (d) zoomed-in view of plot (a) with $\eta=20$, and (e) zoomed-in view near the extremal limit $M=\sqrt{Q^{2}+\alpha}$.
@@ V.3 Insights from the curvature scalar
The expression for the thermodynamic curvature $R_{C}$ obtained from Eq. (32) is lengthy, therefore we refrain from writing it down here; we instead summarize the features extracted from Fig. 6 for representative values of $(\eta,\alpha)$. A robust property across all panels is that $R_{C}\to 0$ for large black hole sizes (large $M$), essentially independent of $(\eta,\alpha)$. In the Ruppeiner interpretation this corresponds to a flat, ideal-gas-like (non-interacting) thermodynamic sector. Such Ruppeiner flatness is known for charged black holes in Einstein gravity across all spacetime dimensions Aman et al. (2003); Aman and Pidokrajt (2006); Masood A. S. Bukhari et al. (2023), and its persistence here provides a direct GR baseline: deviations appear only in the small-horizon/near-extremal regime where $(\alpha,\eta)$ become relevant.
@@
The extremal endpoint also corresponds to $T_{BH}\to 0$ [Fig. 2] and hence to a remnant-like termination of the evaporation picture in our fixed-$Q$ framework. Since $C_{Q}$ and $R_{C}$ become unphysical beyond extremality in this setup, we do not extend the analysis past the endpoint. Nevertheless, it is suggestive to compare with ordinary low-temperature thermodynamics: as $T_{BH}\to 0$, interactions can freeze out, leaving behind quantum-statistical interactions as in ideal Fermi or Bose gases Ruppeiner and Sturzu (2023). In our case, the divergence $R_{C}\to$-$\infty$ at $T_{B}=0$ may therefore be read heuristically as an extreme attractive-correlation signature compatible with a frozen, remnant-like configuration.
A microscopic explanation of these features remains challenging. Black hole thermodynamics is itself subtle at the microscopic level Davies (1978); Page (2005), and the simultaneous presence of higher-curvature corrections (through $\alpha$) and short-distance entropy deformations (through $\eta$) complicates a direct microscopic interpretation. We may, however, still infer certain qualitative features from the present ansatz. From Eq. 15 (or Eq. 13), the horizon scale is primarily controlled by $M$, so the influence of $\alpha$ becomes pronounced only at small radii (remember $\alpha$ is tightly constrained phenomenologically Fernandes et al. (2022)). In this sense, $\alpha$ behaves in a $Q$-like manner in shrinking the horizon scale relative to the GR limit, thereby shifting the near-extremal regime where the thermodynamic-geometry signatures become nontrivial. The macroscopic Ruppeiner flatness may be viewed as reflecting an effective cancellation of attractive and repulsive interaction signatures, while at quantum scales one sign can dominate; in our plots the near-extremal sector is consistently dominated by the attractive ($R_{C}<0$) signature.
Finally, we emphasize that standard Ruppeiner geometry is rooted in equilibrium thermodynamic fluctuations, whereas at quantum scales non-equilibrium ingredients may become relevant Pourhassan et al. (2021b, 2022c); Iso and Okazawa (2011); Pourhassan et al. (2021a). Here we incorporate quantum effects through an entropy deformation without modifying the spacetime geometry, so the curvature analysis should be interpreted as an effective thermodynamic diagnostic based on a quantum-corrected entropy rather than a fully microscopic derivation. For related discussions of repulsive/attractive microstructure interpretations using Ruppeiner geometry in Gauss–Bonnet gravity, see, e.g., Wei and Liu (2020); Wei et al. (2019).
@@ V.4 Theoretical and observational prospects
In the early-Universe context, near-extremal endpoints with $T_{BH}\to 0$ are relevant for primordial black holes (PBHs) Carr and Hawking (1974). In semiclassical GR, sufficiently light PBHs would evaporate completely, whereas an evaporation endpoint at finite mass would instead leave long-lived relics Carr et al. (2021). Such PBH relic/remnant scenarios have been discussed as potential contributors to (or constituents of) dark matter Dalianis and Tringas (2019); Carr and Kuhnel (2022) and are constrained by a broad set of cosmological and astrophysical bounds across PBH mass ranges Carr et al. (2021). In our effective setup, the 4D-EGB background shifts the extremality bound from the GR/RN value to $M=\sqrt{Q^{2}+\alpha}$, implying a modified terminal mass scale (and thus a modified remnant endpoint) in a fixed-$Q$ evaporation picture. This suggests that, if a similar shift persists in a fully dynamical treatment (including charge loss and backreaction), higher-curvature effects could alter the parameter space in which PBH remnants are produced and survive Sasaki and Wang (2025). A natural next step is therefore to embed the present thermodynamic diagnostics into an early-Universe PBH evolution model and juxtapose the shifted-endpoint scenario with existing PBH abundance constraints Carr et al. (2021); Carr and Kuhnel (2022).
@@
Beyond these cosmological implications, $\alpha$ also interfaces with strong-field tests of gravity in the compact-object sector. In particular, it is useful to treat $\alpha$ as an effective strong-field parameter whose physical interpretation (and hence constraining process) is tied up with specific four-dimensional completion/regularization adopted for the underlying 4D-EGB construction Fernandes et al. (2022). A concrete motivation for focusing on strong-field probes is that, in 4D-EGB black-hole spacetimes, quasinormal-mode spectra and photon-sphere/shadow observables acquire explicit $\alpha$-dependence, while stability considerations can already impose nontrivial restrictions on the admissible range of $\alpha$ (see Konoplya and Zinhailo (2020) and references therein). These are precisely the sectors that connect most directly to current and near-future observational programs, including gravitational-wave ringdown measurements Abbott et al. (2016a) and horizon-scale imaging by the Event Horizon Telescope Akiyama et al. (2019). In parallel, astrophysical consistency studies performed in scalar–tensor realizations motivated by higher-dimensional GB gravity indicate that demanding viable compact objects across masses and radii can yield stringent, yet model-dependent, restrictions on the effective coupling Charmousis et al. (2022). Within this landscape, our thermodynamic analysis plays a complementary role: rather than attempting to extract bounds directly, it highlights the region of state space where thermodynamic response functions and information-geometric signatures become maximally sensitive to $\alpha$, thereby indicating where a fully dynamical strong-field confrontation with data (including backreaction, charge loss, and/or quantum-corrected geometry) would be most informative.
@@ VI Conclusion and Outlook
In this work we investigated the thermodynamic phase structure of a charged four-dimensional Einstein–Gauss–Bonnet (4D-EGB) black hole in the GR branch, incorporating quantum-gravity-motivated non-perturbative (exponential) corrections to the entropy while keeping the spacetime geometry, and hence the Hawking temperature defined by surface gravity, classical. Within this effective setup (fixed $Q$ canonical picture), we combined standard thermodynamic diagnostics (heat capacity and Helmholtz free energy) with two complementary probes: (i) an effective quantum-work functional derived from the free-energy landscape via Jarzynski equality, and (ii) information-geometric diagnostics based on the Ruppeiner scalar curvature.
Our results indicate that in the macroscopic regime, the response functions approach the corresponding general-relativistic behavior, consistent with the expectation that both Gauss–Bonnet (GB) and entropy-deformation effects ($\eta$) are suppressed at large horizon scales. In contrast, in the small-horizon/near-extremal regime—where the exponential entropy deformation becomes relevant—the interplay between the GB coupling ($\alpha$) and $\eta$ can generate additional thermodynamic instabilities and curvature singularities. In particular, the endpoint of the fixed-$Q$ evolution is controlled by the extremal limit $M=\sqrt{Q^{2}+\alpha}$, which is shifted relative to the RN baseline ($\alpha=0$) and thereby modifies the inferred terminal mass scale and the associated stability diagnostics in a parameter-dependent manner. The information-geometric analysis provides an internally consistent cross-check: divergences of the Ruppeiner curvature track the loci of thermodynamic singularities, while the sign of the curvature offers a qualitative diagnostic of effective correlations in the thermodynamic state space. At the same time, since quantum effects are implemented here through an entropy deformation on a classical background, any microscopic interpretation should be regarded as model-dependent.
@@
Several extensions are natural. It would be of interest to generalize the analysis to rotating solutions and to include a negative cosmological constant in an extended thermodynamic framework, where additional phase structure and Hawking-Page-type phenomena may arise Kubiznak et al. (2017); Hawking and Page (1983). More fundamentally, a fully dynamical assessment would require incorporating quantum-corrected geometry and backreaction effects to test the robustness of the near-extremal behavior and remnants reported in this study.

### v5

@@ V.2 Computing the Ruppeiner thermodynamic curvature $R_{C}$
[…]
In this framework, the following interpretation is usually ascribed to $R_{C}$: a zero curvature indicates that the system is non-interacting and in an ideal gas-like configuration, providing no additional information about black hole micromolecules. A positive $R$ is associated with repulsive effective interactions, while negative $R$ indicates attractive interactions. Divergences in $R$ correspond to phase transitions [106]. From this interpretation of $R$, one might wonder if $R$ would be very large for a black hole due to its infinite density. This point is emphasized in Ref. [106], where it is suggested that the gravitational degrees of freedom in a black hole system might possess some non-statistical description because all the gravitating material has been compressed into the central singularity. Thus, thermodynamic curvature represents some kind of non-gravitational interactions among the black hole constituents at its surface, arising effectively from the underlying gravitational degrees of freedom.
Figure 6: Thermodynamic curvature $R_{C}$ versus black hole size: impact of (a) quantum corrections $\eta$, (b) positive GB coupling parameter $\alpha$, and (c) negative GB coupling parameter $\alpha$; (d) zoomed-in view of plot (a) with $\eta=20$, and (e) zoomed-in view near the extremal limit $M=\sqrt{Q^{2}+\alpha}$.
@@ V.3 Insights from the curvature scalar
The expression for the thermodynamic curvature $R_{C}$ obtained from Eq. (32) is lengthy, therefore we refrain from writing it down here; we instead summarize the features extracted from Fig. 6 for representative values of $(\eta,\alpha)$. A robust property across all panels is that $R_{C}\to 0$ for large black hole sizes (large $M$), essentially independent of $(\eta,\alpha)$. In the Ruppeiner interpretation this corresponds to a flat, ideal-gas-like (non-interacting) thermodynamic sector. Such Ruppeiner flatness is known for charged black holes in Einstein gravity across all spacetime dimensions [5, 6, 85], and its persistence here provides a direct GR baseline: deviations appear only in the small-horizon/near-extremal regime where $(\alpha,\eta)$ become relevant.
@@
The extremal endpoint also corresponds to $T_{BH}\to 0$ [Fig. 2] and hence to a remnant-like termination of the evaporation picture in our fixed-$Q$ framework. Since $C_{Q}$ and $R_{C}$ become unphysical beyond extremality in this setup, we do not extend the analysis past the endpoint. Nevertheless, it is suggestive to compare with ordinary low-temperature thermodynamics: as $T_{BH}\to 0$, interactions can freeze out, leaving behind quantum-statistical interactions as in ideal Fermi or Bose gases [102]. In our case, the divergence $R_{C}\to$-$\infty$ at $T_{B}=0$ may therefore be read heuristically as an extreme attractive-correlation signature compatible with a frozen, remnant-like configuration.
A microscopic explanation of these features remains challenging. Black hole thermodynamics is itself subtle at the microscopic level [30, 88], and the simultaneous presence of higher-curvature corrections (through $\alpha$) and short-distance entropy deformations (through $\eta$) complicates a direct microscopic interpretation. We may, however, still infer certain qualitative features from the present ansatz. From Eq. 15 (or Eq. 13), the horizon scale is primarily controlled by $M$, so the influence of $\alpha$ becomes pronounced only at small radii (remember $\alpha$ is tightly constrained phenomenologically [34]). In this sense, $\alpha$ behaves in a $Q$-like manner in shrinking the horizon scale relative to the GR limit, thereby shifting the near-extremal regime where the thermodynamic-geometry signatures become nontrivial. The macroscopic Ruppeiner flatness may be viewed as reflecting an effective cancellation of attractive and repulsive interaction signatures, while at quantum scales one sign can dominate; in our plots the near-extremal sector is consistently dominated by the attractive ($R_{C}<0$) signature.
Finally, we emphasize that standard Ruppeiner geometry is rooted in equilibrium thermodynamic fluctuations, whereas at quantum scales non-equilibrium ingredients may become relevant [97, 96, 62, 93]. Here we incorporate quantum effects through an entropy deformation without modifying the spacetime geometry, so the curvature analysis should be interpreted as an effective thermodynamic diagnostic based on a quantum-corrected entropy rather than a fully microscopic derivation. For related discussions of repulsive/attractive microstructure interpretations using Ruppeiner geometry in GB gravity, see, e.g., refs. [120, 119].
@@ V.4 Theoretical and observational prospects
In the early-Universe context, near-extremal endpoints with $T_{BH}\to 0$ are relevant for primordial black holes (PBHs) [19]. In semiclassical GR, sufficiently light PBHs would evaporate completely, whereas an evaporation endpoint at finite mass would instead leave long-lived relics [20]. Such PBH relic/remnant scenarios have been discussed as potential contributors to (or constituents of) dark matter [29, 21] and are constrained by a broad set of cosmological and astrophysical bounds across PBH mass ranges [20]. In our effective setup, the 4D-EGB background shifts the extremality bound from the GR/RN value to $M=\sqrt{Q^{2}+\alpha}$, implying a modified terminal mass scale (and thus a modified remnant endpoint) in a fixed-$Q$ evaporation picture. This suggests that, if a similar shift persists in a fully dynamical treatment (including charge loss and backreaction), higher-curvature effects could alter the parameter space in which PBH remnants are produced and survive [109]. A natural next step is therefore to embed the present thermodynamic diagnostics into an early-Universe PBH evolution model and juxtapose the shifted-endpoint scenario with existing PBH abundance constraints [20, 21].
@@
Beyond these cosmological implications, $\alpha$ also interfaces with strong-field tests of gravity in the compact-object sector. In particular, it is useful to treat $\alpha$ as an effective strong-field parameter whose physical interpretation (and hence constraining process) is tied up with specific four-dimensional completion/regularization adopted for the underlying 4D-EGB construction [34]. A concrete motivation for focusing on strong-field probes is that, in 4D-EGB black hole spacetimes, quasinormal-mode spectra and photon-sphere/shadow observables acquire explicit $\alpha$-dependence, while stability considerations can already impose nontrivial restrictions on the admissible range of $\alpha$ (see ref. [68] and references therein). These are precisely the sectors that connect most directly to current and near-future observational programs, including gravitational-wave ringdown measurements [1] and horizon-scale imaging by the Event Horizon Telescope [4]. In parallel, astrophysical consistency studies performed in scalar–tensor realizations motivated by higher-dimensional GB gravity indicate that demanding viable compact objects across masses and radii can yield stringent, yet model-dependent, restrictions on the effective coupling [24]. Within this landscape, our thermodynamic analysis plays a complementary role: rather than attempting to extract bounds directly, it highlights the region of state space where thermodynamic response functions and information-geometric signatures become maximally sensitive to $\alpha$, thereby indicating where a fully dynamical strong-field confrontation with data (including backreaction, charge loss, and/or quantum-corrected geometry) would be most informative.
@@ VI Conclusion and Outlook
In this work, we investigated the thermodynamic phase structure of a charged 4D-EGB black hole in the GR branch, incorporating quantum-gravity-motivated non-perturbative (exponential) corrections to the entropy while keeping the spacetime geometry, and hence the Hawking temperature defined by surface gravity, classical. Within this effective setup (fixed $Q$ canonical picture), we combined standard thermodynamic diagnostics (heat capacity and Helmholtz free energy) with two complementary probes: (i) an effective quantum-work functional derived from the free-energy landscape via Jarzynski equality, and (ii) information-geometric diagnostics based on the Ruppeiner scalar curvature.
Our results indicate that in the macroscopic regime, the response functions approach the corresponding general-relativistic behavior, consistent with the expectation that both GB and entropy-deformation effects ($\eta$) are suppressed at large horizon scales. In contrast, in the small-horizon/near-extremal regime—where the exponential entropy deformation becomes relevant—the interplay between the GB coupling ($\alpha$) and $\eta$ can generate additional thermodynamic instabilities and curvature singularities. In particular, the endpoint of the fixed-$Q$ evolution is controlled by the extremal limit $M=\sqrt{Q^{2}+\alpha}$, which is shifted relative to the RN baseline ($\alpha=0$) and thereby modifies the inferred terminal mass scale and the associated stability diagnostics in a parameter-dependent manner. The information-geometric analysis provides an internally consistent cross-check: divergences of the Ruppeiner curvature track the loci of thermodynamic singularities, while the sign of the curvature offers a qualitative diagnostic of effective correlations in the thermodynamic state space. At the same time, since quantum effects are implemented here through an entropy deformation on a classical background, any microscopic interpretation should be regarded as model-dependent.
@@
Several extensions are natural. It would be of interest to generalize the analysis to rotating solutions and to include a negative cosmological constant in an extended thermodynamic framework, where additional phase structure and Hawking-Page-type phenomena may arise [70, 48]. More fundamentally, a fully dynamical assessment would require incorporating quantum-corrected geometry and backreaction effects to test the robustness of the near-extremal behavior and remnants reported in this study.